OpenStudy (anonymous):

2x over 3 +6=12 solve?

OpenStudy (anonymous):

\[\frac{2x}{3}+6=12\]?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

solve for x?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

subtract 6 from both sides 2x/3+6-6=12-6 2x/3=6 multiply both sides with 3 2x/2*(3)=6*3 2x=18 divide both sides by 2 2x/2=18/2 x=9

OpenStudy (anonymous):

first subtract 6 from both sides \[\frac{2x}{3}+6-6=12-6\] \[\frac{2x}{3}=6\] multiply by 3 to get rid of the fraction \[3(\frac{2x}{3})=6(3)\] \[2x=18\] divide both sides by 2 \[\frac{2x}{2}=\frac{18}{2}\] the 2/2 is 1 \[x=9\]

OpenStudy (anonymous):

oh okay thanks guys

OpenStudy (anonymous):

yes that the answer.

OpenStudy (anonymous):

am getting x = 15 when i begin with 12*3

OpenStudy (anonymous):

2x/3+6=12 2x+6=12*3 X=(36-6)/2 X=15